OpenStudy (goformit100):

Prove that there are no natural numbers, which are solutions of 15x^2 - 7y^2 =9

OpenStudy (goformit100):

@funinabox

OpenStudy (zarkon):

I would start by showing that x and y are both multiples of 3

OpenStudy (goformit100):

ok , what to do after that ?

OpenStudy (anonymous):

Re-arrange:\[\bf y^2=\frac{ 15x^2-9 }{ 7 }=\frac{ 3 }{ 7 } (5x^2-3)\]If the solutions are natural numbers, then they must be non-negative, hence:\[\bf y=+\frac{ \sqrt{3} }{ \sqrt{7} }\sqrt{5x^2-3}\]This can be re-written as:\[\bf \frac{ y }{ \sqrt{5x^2-3}}=\frac{ \sqrt{3} }{ \sqrt{7} }\]We must note that sqrt{3} and sqrt{7} are both irrational. This implies that 'y' must also be irrational. Comparing the denominators we get:\[\bf 5x^2-3=7 \implies x = \pm \sqrt{2}\], and sqrt{2} is also irrational. Hence there is no natural number values of "x" and "y" which satisfy the given equation.

OpenStudy (anonymous):

@goformit100

OpenStudy (goformit100):

Thank you sir

OpenStudy (zarkon):

@genius12 are you saying that both x and y have to be irrational to satisfy the original equation?